Willingness to Pa S0 $13 $22 $27 $33 $60 Bu Dan Aria Hanna Emi Caleb 10. The table above refers to six possible buyers willingness to pay to watch a clown perform Assume these are the only buyers in the market, and they are rational. Based on the information in the table, if the tickets to the clown show sell for $20 each, then consumer surplus in this market will be: a. Infinitely negative b. 35 C. 62 d. 135 e. 155
Answer #1

Answer : 10) The answer is option c.

Because if the market price level is $20 for each "clown show" ticket then Dan and Spencer will not buy the tickets. Because Dan is ready to pay $0 and Spencer is ready to pay only $13 for each ticket. So, except Dan and Spencer other 4 buyers will buy the tickets.

Now,

Consumer surplus for Aria = Willingness to pay - market price level = 22 - 20 = $2.

Consumer surplus for Hanna = Willingness to pay - market price level = 27 - 20 = $7.

Consumer surplus for Emily = 33 - 20 = $13

Consumer surplus for Caleb = 60 - 20 = $40.

Therefore, the market consumer surplus = 2 + 7 + 13 + 40 = $62.

Thus, option c is the correct answer.